Session: Feb 2005 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Tao An at Thu Feb 17 15:24:51 2005) Success :

Medicina (Andrea Orlati at Thu Feb 17 16:00:30 2005) Success :
light rain during first 2 hours of the observation

Urumqi (Liu Xiang at Fri Feb 18 02:27:55 2005) Success :
It is very clear, with little part of snow in the antenna surface. 8:00--8:14 antenna pc problem. disk2file didn't produce file automatically, we will manually retrieve 2 small scans and ftp to Bologna.

Onsala (Michael Lindqvist at Fri Feb 18 07:29:39 2005) Success :
No known problems.

Effelsberg (A. Kraus at Fri Feb 18 07:47:18 2005) Minor failures : TSYS SLEW PCAL
Some problems in the beginning: no Phasecal, no proper Tsys. Reboot of FS at 1040UT, back on source at UT1152, everything fine. Slewing problems (too less time) from UT 1508 on: fixed antenna on 0133+476.

Noto (P. Cassaro at Sat Feb 19 15:36:55 2005) Minor failures :
The schedule refused to create the ftp files.

Metsahovi (Ari Mujunen at Mon Feb 21 13:09:17 2005) Minor failures : LATE
Four first minutes lost because a large azeloff was left in antenna control after pointing observations. New phase cal system on, injected to IF in receiver in antenna.

Cambridge (Alastair Gunn at Mon Feb 21 20:58:20 2005) Success :
No known problems.

Jodrell_Bank (Alastair Gunn at Mon Feb 21 20:59:07 2005) Success :
MkII was used. No known problems.